OpenStudy (anonymous):

x^2y-2x+3y=6 differentiation to find tangent on a line help plzz

OpenStudy (dumbcow):

so you are trying to solve for dy/dx or y' ??

OpenStudy (anonymous):

implicit

OpenStudy (anonymous):

trying to find the tangent of the line but u have find d/dx first

OpenStudy (dumbcow):

\[\frac{dy}{dx} = \frac{2-2xy}{x^{2}+3}\]

OpenStudy (anonymous):

y? i dnt understand hw u get that

OpenStudy (dumbcow):

by using implicit differentiation, you take derivative of each term but when you have a "y" the derivative is dy/dx \[\rightarrow 2xy + x^{2}\frac{dy}{dx} -2 +3\frac{dy}{dx} = 0\] then isolate and solve for dy/dx

OpenStudy (anonymous):

ohhhh